Let X = {0, 1, 2}. List the elements of each of the following sets:
a. {z : z = 2x and x ∈ X}
b. {z : z = x + y where x ∈ X and y ∈ X}
1
Expert's answer
2020-11-17T07:04:30-0500

a: let's iterate over the elements in the set x to make new set

"2*0=0" , "2*1=2" , "2*2=4"

it means set is {0, 2, 4}

b: let's iterate over the elements in the set x to make new set

"0+0=0, 0+1=1,0+2=2,"

"1+1=2,1+2=3,2+2=4"

it means set is {0, 1, 2, 3, 4}